OpenStudy (anonymous):

What is the second derivative of y= 8x^(1/3). Possible answers are: 48 16 -32 32 -16 I am quite confused on this concept of derivative but this is what I have: d^2/dx^2*f(x)= 8x^(1/3) fx= d^2/dx^2 * 8x^(1/3) fx= d/dx * (d/dx* 8x^(1/3) fx= d/dx * (8*d/dx*x^(1/3) fx= d/dx * (8*1/3*x^(1/3-1) fx= d/dx * (8/3*x^(-2/3) fx=8/3* d/dx * x^(-2/3) fx=8/3*-2/3* x^(-2/3-1) fx=8/3*-2/3* x^(-2/3-1) fx=-16/9* x^(-5/3) fx=-16/(9* x^(-5/3)) But how do I get from here to my answers??

OpenStudy (anonymous):

y=8*x^(1/3). First derivative: y(') = 8 * (1/3) * x^(1/3 - 1) = (8/3) * x^(-2/3) Second derivative: y('')=(8/3)*(-2/3)*x^(-2/3 - 1)= (-16/9) * x^(-5/3). Or, to put it in a different form y('')= -16/ ( 9 * x ^ (5/3) )= -16 / (9* cuberoot (x^5) ) So, you're spot on with the derivative imo. The only thing that's missing here is a value. Are you sure there's nothing missing from the problem ? A context of some sort ? Or maybe a typo like x^3 instead of x^ (1/3).

OpenStudy (anonymous):

"Find the second derivative" is the only context.